OpenStudy (anonymous):

The legend of a map indicates that 1 inch=16 miles. If the distance on the map between two cities is \[2\frac38\] inches, how far apart are the cities?

OpenStudy (anonymous):

Make the fraction an improper fraction then multiply it by 16miles.

OpenStudy (anonymous):

so the improper fraction is \[\frac{19}{8}\times \frac{16}{1}=\]

OpenStudy (anonymous):

how do the miles and inches cancel out?

OpenStudy (anonymous):

\[{16 \over 1}=16\]

OpenStudy (anonymous):

Do you understand? :)

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

Good. :) now just multiply \[{19 \over 8} \times 16\]

OpenStudy (anonymous):

You can solve it now, I believe?

OpenStudy (anonymous):

I can do the multiplications, and the answer is 38. i just don't understand how the inches cancel out to give miles

OpenStudy (anonymous):

Here's an example: You have 2 inches and 1 inch on the map equals 10miles. You would then proceed to multiply 2 by 10, right? this also can be written as 2/1 times 10. It really is just as simple as that. (I wanted to type a longer expiation but after thinking and thinking, i thought this would be suffice.)

OpenStudy (anonymous):

explanation*

OpenStudy (anonymous):

Unless the question states that 2,3,4 inches etc., you don't have to worry about converting the miles proportioned to the inches.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

you're welcome. I hope you understand. If not, hopefully a little; and as you keep doing questions like this, you will begin to understand why.